OpenStudy (anonymous):

deduce that \[\lim_x\rightarrow\infty{\frac{x^2}{n!}}\] = 0 for all x

OpenStudy (anonymous):

the question is wrong... here it is again:\[\text{Deduce that }\lim_{n\rightarrow\infty}{\frac{x^2}{n!} = 0}\text{ for all x}\]

OpenStudy (anonymous):

it's still wrong, it must be x^n and not x^2 :-( I'm sleepy

OpenStudy (zarkon):

ok...that better

OpenStudy (zarkon):

show that the series \[\sum_{n=1}^{\infty}\frac{x^n}{n!}\] has radius of convergence of infinity. then the limit of the sequence has to be zero

OpenStudy (anonymous):

I have shown it, so do I just say that 'since that converges for all x, then for all x the limit of the sequence must be 0'

OpenStudy (zarkon):

if it wasn't zero for a specific value of x then the series would not converge by the divergence test

OpenStudy (anonymous):

right

OpenStudy (anonymous):

alright I'm done with math. \begin{array}l\color{#FF0000}{\text{Y}}\color{#FF0000}{\text{ }}\color{#FF7F00}{\text{i}}\color{#FF7F00}{\text{ }}\color{#FFE600}{\text{p}}\color{#FFE600}{\text{ }}\color{#00FF00}{\text{p}}\color{#00FF00}{\text{ }}\color{#0000FF}{\text{e}}\color{#0000FF}{\text{ }}\color{#6600FF}{\text{e}}\color{#6600FF}{\text{ }}\color{#8B00FF}{\text{!}}\color{#8B00FF}{\text{ }}\color{#FF0000}{\text{!}}\color{#FF0000}{\text{ }}\color{#FF7F00}{\text{!}}\color{#FF7F00}{\text{ }}\end{array} no just kidding I will do 6 more chapters again tomorrow :-P

OpenStudy (zarkon):

nice

OpenStudy (anonymous):

My fault for missing 3 weeks of school

OpenStudy (zarkon):

that's not good

OpenStudy (anonymous):

hope you had fun at least!